This is Graphing Linear Equations. Can you please help me solve this problem?
x+2y=-8. Thank you for any help you can offer.

Answer by checkley71(8403)

THE STANDARD FORMULA FOR A LINE IS Y=mX+b WHEN m=THE SLOPE & b=THE Y INTERCEPT
X=2Y=-8
2Y=-X-8
2Y=-X-8
Y=-X/2-8/2
Y=-X/2-4
THEN THE SLOPE=-1/2 AND THE Y INTERCEPT=-4
YOU CAN ALSO FIND THE X INTERCEPT BYU SETTING Y=0 AND SOLVIONG FOR X. THUS:
0=-X/2-4
-X/2=4
X=-2*4
X=-8 OR (-8,0)& THE Y INTERCCEPT IS EXPRESSED AS (0,-4)